a vessel contains CO2 and CO with pressures 2 atm and 3 atm respectively at 27’c . at a temperature of 2727’C , thje reaction 2CO2 = CO2 + C occurs and equilibrium is attained . if the equilibrium pressure is 45 atm , then what is the kp of the reaction ?

Grade:12th pass

2 Answers

Arun
5 years ago
Dear student
 
I think there is some mistake in question. Because I am getting sime unusual answer.
Perhaps pressure should be 4.5 atm.
 
Please check.
 
Regards
Arun (askIITians forum expert)
Harshit Rana
4 years ago
 Calculate the pressure of Co2 and Co each at 2727•C Using P1/T1 =P2/T2  ,Pco =20atm ,Pco2=30atm    At Equilibrium Pco =30-2P  ,  Pco2 =20+P +30-2P+20+P=45  ,   P= 5atm %dissociation of co =10/30x100 =33.33%